OpenStudy (anonymous):

The area of a rectangular wall of a barn is 140 square feet. Its length is 6 feet longer than twice its width. Find the length and width of the wall of the barn.

OpenStudy (aravindg):

its too easy! have you really tried this qn? If yes tell me where you are stuck?

OpenStudy (anonymous):

the width is is 3ft stuck on the length

OpenStudy (anonymous):

@joanna601 Start with writing down the formula of Area on HERE ....

OpenStudy (anonymous):

area=lw

OpenStudy (anonymous):

Do you know how to express :"twice its width"

OpenStudy (anonymous):

2w

OpenStudy (anonymous):

Good, continue with "6 feet longer than twice its width"

OpenStudy (anonymous):

l=6+2w

OpenStudy (anonymous):

Great :) Now you plug into: A = L W .. = ....

OpenStudy (anonymous):

I know the width is 3 but I am lost on length

OpenStudy (anonymous):

Haven't you just answered me L = 2W + 6 ???

OpenStudy (anonymous):

A = W * L = W ( 2W + 6) = 140 Can you continue?

OpenStudy (anonymous):

why is there a blank in front of the equation?